A family of 2 adults and 5 children have their names in a bowl. There needs to be picked 3 names. I have to find the probability of it being 1 adult and 2 children
Answer by ewatrrr(24785)


Hi
Family of 7:  2A  5C
Pick three names from bowel

P(1A & 2C) =  (2C1)(5C2)/(7C3) = 20/35 = 4/7

Wish You the Best in your Studies.